OpenStudy (alexandrad):

Write the equation 4x + 2y = 5 in the form y = mx + b.

OpenStudy (steve816):

Hey there. Here is a hint to get you started: Subtract 4x from both sides!

OpenStudy (alexandrad):

A. y=2x+5/2 B. y=2x+5 C. y=-2x+5/2 D. y=-2x-5/2

OpenStudy (alexandrad):

so 4x-4= o and 5-4=1 so 2y=1

OpenStudy (alexandrad):

i divide by 2 and get 2 so y=2

OpenStudy (steve816):

Not quite. Here is what you do! You have to first subtract 4x from both sides.\[\Huge 2y = -4x +5\]

OpenStudy (steve816):

Now, you want to get the y alone. So divide by 2 to both sides, and the answer would be:\[\Huge y = -2x + \frac{ 5 }{ 2 }\]

OpenStudy (alexandrad):

thank you for explaining this to me

OpenStudy (steve816):

No problem Alexandra!